OpenStudy (anonymous):

Solve for x. x/5 - x/6 = 1/3 x = 1/10 x = 1/90 x = 10

OpenStudy (zepp):

What you would get if you try to put \(\large \frac{x}{5} and \frac{x}{6}\) under a same denominator?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

x=10

OpenStudy (zepp):

E.g.: \(\huge \frac{1}{2} - \frac{1}{3} = \frac{3}{2*3}-\frac{2}{2*3}\)

OpenStudy (diyadiya):

@studyisbest Please don't give the final Answer! Its against the rules of Openstudy

OpenStudy (zepp):

Thank you for cutting me off, @studyisbest

OpenStudy (diyadiya):

@studyisbest http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok tahnks for info @Diyadiya